OpenStudy (anonymous):

Please help. (9/3x-1)-(5x/2x+3)

OpenStudy (anonymous):

\[\frac{9}{3x-1}-\frac{5x}{2x+3}\] is my guess

OpenStudy (mathmale):

I'd guess the main reason why you haven't gotten any responses until now is that your expression is a bit hard to decipher. I suspect your (9/3x-1) should be 9 / (3x-1). Likewise, (5x/2x+3) should possibly have been (5x) / (2x+3). I strongly recommend that you learn to use Equation Editor, or use the Draw feature. Note how satellite73 says, "is my guess"? He, too, has had to guess. Clarity in presenting your math expressions is very important.

OpenStudy (anonymous):

except i wrote it wrong it is \[\large\frac{A}{B}-\frac{C}{D}=\frac{AD-BC}{BD}\]

OpenStudy (mathmale):

@surfergirl14 Please help. (9/3x-1)-(5x/2x+3) The rest is up to you. Please clarify what you meant. Use the draw utility, OK? If you don't want to, then use more parentheses and place them appropriately.

OpenStudy (anonymous):

\[9/(3x-1) - (5x)/(2x+3)\] is that better? I don't know that's all the book gave me

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

thank you. That's SO much clearer, so much better.

OpenStudy (anonymous):

ok good. yea i'm terrible at math, I even went back over the lesson and just can't get it.

OpenStudy (mathmale):

One favor I need to ask you: Please stop describing yourself as "terrible in math." It does not help you in any way. Look at math as a challenge that you can overcome.

OpenStudy (mathmale):

\[9/(3x-1) - (5x)/(2x+3)\rightarrow \frac{ 9 }{ 3x-1 }-\frac{ 5x }{2x+3 }\]

OpenStudy (mathmale):

I'm assuming you're familiar with the concept of "lowest common denominator." Can you identify the LCD in this problem?

OpenStudy (anonymous):

how can I get an LCD when they're equations or not like for example 12 or 8? get what i'm saying?

OpenStudy (mathmale):

\[In~\frac{ 1 }{ 3}-\frac{ 1 }{ 8 },\] the LCD is found by multiplying the two different denominators together: (3)(8)=24. Sound familiar? If so, finish combining these terms.

OpenStudy (mathmale):

To answer your question: Your task here is to combine two rational fractions into one. It makes no difference whether you have an expression (as you do here) or an equation.

OpenStudy (mathmale):

\[\frac{ 1 }{ 3 }-\frac{ 1 }{ 8 }=\frac{ ? - ? }{ 24 }\]

OpenStudy (mathmale):

Ms. Surfer?

OpenStudy (anonymous):

sorry my wifi is being weird